OpenStudy (anonymous):

A circle whose center is at(6,8) passes through the origin. Which of the following points is NOT on the circle?

ganeshie8 (ganeshie8):

where are the following points, you got options ?

OpenStudy (anonymous):

A. (12,0), B(6,-2), C(16,8), D(-2,12), E(-4,8)

OpenStudy (phi):

I would first write down the equation for this circle. you know its center, and you can find its radius then use (x-x0)^2 + (y-y0)^2 = r^2 then test each of the points in the equation

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

distance formula

OpenStudy (anonymous):

length is 10

OpenStudy (anonymous):

for radius

OpenStudy (anonymous):

so diameter is 20

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

6-8-10 right triangle?

OpenStudy (anonymous):

I was able to make a 6-8-10 right triangle

ganeshie8 (ganeshie8):

wait

ganeshie8 (ganeshie8):

you got radius = 10

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

I got points (0,0),(6,0), and (6,8)

OpenStudy (phi):

I would just use (x-6)^2 + (y-8)^2 = 100 one of the pair will not work

ganeshie8 (ganeshie8):

just check distance between center and each of the given points

OpenStudy (anonymous):

ahh

ganeshie8 (ganeshie8):

if u get 10, then that point lies on circle.

OpenStudy (anonymous):

so like

OpenStudy (anonymous):

ahh

OpenStudy (phi):

that is what the equation of a circle does

OpenStudy (anonymous):

D. does not work

OpenStudy (anonymous):

square root of 148 is an incorrect answer

OpenStudy (anonymous):

ty